Use the rational zeros theorem to find all the real zeros of the polynomial function. Use the zeros to factor f over the real numbers.
f(x)=x^3-2x^2-13x-10
Find the real zeros of f.
x=?
Use the real zeros to factor f.
f(x)=?

Answer by lwsshak3(11628)

Use the rational zeros theorem to find all the real zeros of the polynomial function. Use the zeros to factor f over the real numbers.
f(x)=x^3-2x^2-13x-10
Find the real zeros of f.
x=?
Use the real zeros to factor f.
f(x)=?
**
Using rational roots theorem:
....0....|....1.......-2.......-13........-10...........
...-1....|....1......-3.......-10..........0 (-1 is a zero)
f(x)=(x+1)(x^2-3x-10)=(x+1)(x+2)(x-5)
real zeros of f: -2, -1, & 5)
f(x)=(x+1)(x+2)(x-5)